OpenStudy (anonymous):

How Do you write "A is the set of even whole numbers less than 12" in the notation builder?

Parth (parthkohli):

\[\{\mbox{name of the set}:\mbox{condition}\}\]

OpenStudy (anonymous):

can you write it for me? i still don't understand

Parth (parthkohli):

What do you call the set of even whole numbers?

OpenStudy (anonymous):

PartKohli? please help

Parth (parthkohli):

Yes, I am helping ya. :-)

OpenStudy (anonymous):

Great,Thanks a bunch! you're very nice :D

OpenStudy (anonymous):

my problem is i don't know how to write it is it like this: A={x|x<...,x \[\in \] ...} (note that "...." means i don't know what to put in there)

Parth (parthkohli):

\[2\mathbb Z\]is the set of even numbers.

Parth (parthkohli):

Where are you exactly stuck...?

OpenStudy (anonymous):

um,i'm stuck in the "..." part.. what should i write there?

OpenStudy (aravindg):

there you should write where the element belongs to

OpenStudy (aravindg):

and also include any condition if present

Parth (parthkohli):

\(x\) is more than or equal to \(0\) and less than or equal to \(12\).

Parth (parthkohli):

For example, you can write the set of first ten natural numbers as:\[\{x\in\mathbb N: x \le 10\}\]

OpenStudy (anonymous):

i see! Tx you two! you guys are very nice for a newcomer like me,haha! well,Tx Again!

Parth (parthkohli):

Can you do it now? :-)

OpenStudy (anonymous):

hmm,kind off... well,Tx!

OpenStudy (aravindg):

for example if you are asked to write set of all even numbers you can write it as A={ x:x=2n ;n \(\in\) N} N is set oof all natural numbers

OpenStudy (aravindg):

*of

OpenStudy (aravindg):

yw :)

Parth (parthkohli):

Nah, I'd just use\[A = \{x:x\in2\mathbb Z \}\]And BTW, @AravindG, it should be the set of integers :-)

OpenStudy (aravindg):

I wanted to make it clear ..in schools we are first taught in that fashion ,then we play with it :) @ParthKohli

Parth (parthkohli):

But even numbers belong to the set of integers, not only N.

OpenStudy (aravindg):

*corrected :positive even numbers